OpenStudy (hyroko):

suppose a true approval rating of a particular candidate in the population is 48% If a simple random sample of 3000 people is done, what is the probability that the resulting percent that approve of the person is above 50%?

OpenStudy (hyroko):

@kropot72 care to help me with one more?

OpenStudy (kropot72):

@Hyroko This can be solved by using the normal approximation to the binomial distribution.

OpenStudy (hyroko):

I'm sorry I don't understand either of those suggestions.

OpenStudy (hyroko):

Can someone actually help me figure this out?

OpenStudy (hyroko):

How do I calculate this without knowing the standard deviation?

OpenStudy (dumbcow):

you can calculate the std dev \[\sigma = \sqrt{\frac{P(1-P)}{n}} = \sqrt{\frac{.48(.52)}{3000}}\] find z-score \[z = \frac{p-P}{\sigma} = \frac{.5-.48}{\sigma}\] Use normal table to find \[P(Z >z) = 1-P(Z < z)\] The table will give probability of less than given z-value